Consider two solids, one (A) with a density of 2.79 g/mL, and the other (B) with a density of 7.91 g/mL.

a) If you mix 1.00 mL of each of the two substances together, what is the density of the mixture?..........g/ml

b) If you mix 1.00 g of each of the two substances together, what is the density of the mixture?..........g/ml




1
Expert's answer
2021-01-21T03:42:25-0500

a) if we mix 1 ml of each of the substances, we get 2 ml of the new mixture. m1 = p*V = 2.79*1 = 2.79 grams, the mass of the first substance. Similarly, the mass the second equals 7.91 grams. The total mass is 2.79 + 7.91 = 10.7 grams. Thus, p = m/V = 10.7 / 2 = 5.35 g/ml.

b) analogically, we get 2 grams of the mixture and the total volume of (1/2.79 + 1/7.91) = 0.48 ml; p = m/V = 2/0.48 = 4.66 g/ml.


SOLUTION: a) 5.35 g/ml, b) 4.66 g/ml.